OpenStudy (anonymous):

Julie stands at the top of a tall tower. She drops a bowling ball from the top of the tower, then One second later, she drops a second bowling ball. While the balls are in the air, which of the following is true about their separation? Ignore any effects due to air resistance. A.) Decreases over time B.) Stays the same C.) Increases overtime Explain why...

OpenStudy (anonymous):

Let y1=1/2gt^2 be vertical position of the mass 1 and y2=1/2g(t+1)^2 be vertical position of mass 2. difference between y1 and y2 become, y2-y1=(1/2gt^2+2t+1)-1/2gt^2-=2t+1 It shows that the discrepancy mass 1 and mass 2 would increase.

OpenStudy (anonymous):

So overtime, the separation between the two bowling balls will increase?

OpenStudy (anonymous):

Yes it should be. So overtime we see that the separation would increase linearly.

OpenStudy (anonymous):

At time t=0, the separation is 1. For time t=1, it would be 3 ...

OpenStudy (anonymous):

Ohhh, okay. I kinda get it now! (: Thankkkks!

OpenStudy (anonymous):

Anytime.